# Online estimation of the inverse of the Hessian for stochastic optimization with application to universal stochastic Newton algorithms

### v1

@@ 1 Introduction
where $\theta_{0}^{SG}$ is an arbitrarily chosen initial value and $(\nu_{n})_{n\geq 1}$ is a sequence of positive real numbers decreasing towards 0. These algorithms have been extensively studied, with asymptotic results found by Pelletier, (1998, 2000) and non-asymptotic results, such as uniform bounds of the quadratic mean error, presented by Moulines and Bach, (2011); Gadat and Panloup, (2017); Godichon-Baggioni, (2021) to name a few. To ensure asymptotic efficiency, an additional step consists of considering an averaged version of the estimates (Polyak and Juditsky,, 1992).
@@
We thus obtain a universal estimate of the inverse of the Hessian, and thus with reduced calculus time. To further enhance convergence rate, we also consider its weighted averaged version, as discussed by Mokkadem and Pelletier, (2011); Boyer and Godichon-Baggioni, (2022). We establish the almost sure rates of convergence for the proposed estimates, after making slight modifications. These results remain true for any consistent estimates $\theta_{n}$. Based on this concept, we introduce a universal recursive Newton algorithm and its weighted averaged version. Additionally, we provide their convergence rates and demonstrate the asymptotic efficiency of the weighted averaged estimates.
@@ 2 Framework
where the loss $g(X,\cdot)$ is a convex, twice-differentiable function and $X$ is a random vector of ${\mathbb{R}}^{p}$. We assume that there exists a unique value $\theta\in{\mathbb{R}}^{d}$ such that
@@
(A1) There exists $C>0$ such that for all $h\in\mathbb{R}^{d}$,
@@
These assumptions are very close to those presented in the literature (Pelletier,, 2000; Gadat and Panloup,, 2017; Godichon-Baggioni,, 2019). Assumption (A1) controls the growth of the gradient and guarantees the stability of the estimation process. It ensures that the gradient remains bounded as the estimation progresses. Assumption (A2) ensures that the curvature of $G$ at $\theta$ is well-behaved, allowing the estimation algorithm to reliably exploit the local structure of $G$. This assumption also guarantees that the gradient of $G$ is Lipschitz continuous with a constant $L_{\nabla G}$. This Lipschitz continuity is crucial as convergence results are obtained using a Taylor’s expansion of $G$ up to the second order. Assumption (A3) indicates that the Hessian matrix does not exhibit abrupt changes within a neighborhood around $\theta$. This assumption ensures the stability of the Hessian estimates during the estimation process. Assumption (A4) guarantees that the second-order derivative of $G$ does not exhibit excessive fluctuations. It imposes a bound on the variation of the Hessian matrix, providing further stability to the estimation algorithm. It’s worth noting that Hölder’s inequality leads to
@@
These assumptions, along with the differentiability properties of $g$ and $G$, provide a solid foundation for developing efficient second order methods to solve the minimization problem and obtain reliable estimates of $\theta$.
@@ 3 Estimation of the Hessian inverse
In this section, our focus is solely on estimating the inverse Hessian of function $G$ in $\theta$, denoted as $H^{-1}$ with $H=\nabla^{2}G(\theta)$. Even if our motivation is to estimate $H^{-1}$ for proposing a second-order algorithm, the estimation of $H^{-1}$ can also be valuable for recursively constructing confidence intervals or significance statistical tests for a component of parameter $\theta$ when the parameter $\theta$ is estimated using another asymptotically efficient algorithm like the averaged stochastic gradient algorithm. Indeed, in most cases, the asymptotic variance involving in the central limit theorem, generally depends on matrix $H^{-1}$ and its estimation is then required.
@@
where $\widehat{A}_{0}$ is an arbitrary symmetric positive definite matrix, and $\gamma_{n}=c_{\gamma}n^{-\gamma}$ with $\frac{1}{2}<\gamma<1$ and $c_{\gamma}>0$. It is important to note that $\widehat{A}_{n}$ is symmetric for any $n\geq 1$ due to its construction. However, since $\theta$ is unknown, we need to estimate it. Assuming we have an efficient recursive estimator of $\theta$ (e.g., a stochastic gradient estimator), we can easily derive an estimator of $H^{-1}$ using a plug-in procedure:
@@
This estimator is always symmetric but not necessarily positive definite. To ensure positive definiteness, we introduce a truncation based on the norm of $\nabla_{h}^{2}g(X_{n},\widehat{\theta}_{n-1})$, leading to the following estimator of $H^{-1}$:
$\widehat{A}_{n}=\widehat{A}_{n-1}-\gamma_{n}\left(\widehat{A}_{n-1}\nabla_{h}^{2}g(X_{n},\widehat{\theta}_{n-1})+\nabla_{h}^{2}g(X_{n},\widehat{\theta}_{n-1})\widehat{A}_{n-1}-2I_{d}\right)\mathbf{1}_{\left\{\left\|\nabla^{2}_{h}g(X_{n},\widehat{\theta}_{n-1})\right\|_{op}\leq\beta_{n}\right\}},$ (3)
where $\beta_{n}=c_{\beta}n^{\beta}$ with $\frac{1-\gamma}{q-1}<\beta<\gamma-\frac{1}{2}$ and $0<c_{\gamma}c_{\beta}<\frac{1}{2}$. Additionally, this truncation enables control over the smallest eigenvalue of $\widehat{A}_{n}$, which is useful for studying an estimator of the parameter $\theta$ involving the matrix $\widehat{A}_{n}$. This is particularly important in establishing the consistency of the Stochastic Newton algorithm presented in Section 4.
@@
$A_{n}$ $=A_{n-1}-\gamma_{n}\left(P_{n}Q_{n}^{T}+Q_{n}P_{n}^{T}-2\,I_{d}\right)\mathbf{1}_{\{\left\|Q_{n}\right\|\left\|Z_{n}\right\|\leq\beta_{n}\}}$ (4)
where $A_{0}$ is an arbitrary symmetric and positive definite matrix.
We can observe that in this algorithm the truncation is only based on $\left\|Q_{n}\right\|\left\|Z_{n}\right\|$, and not on $\left\|Q_{n}Z_{n}^{T}\right\|_{op}$ as expected, because we have $\left\|Q_{n}Z_{n}^{T}\right\|_{op}\leq\left\|Q_{n}\right\|\left\|Z_{n}\right\|$. Notably, the computational complexity of each update of $A_{n}$ is now reduced to $\mathcal{O}(d^{2})$. Moreover, following Mokkadem and Pelletier, (2011); Boyer and Godichon-Baggioni, (2022), we can propose a weighted averaged estimate $A_{n,\tau}$, which performs better in practice when the initialization is poor. It is given by
@@
Assume that Assumptions (A2) to (A4) hold, and that there is an estimate $\hat{\theta}_{n}$ satisfying for all $\delta>0$
@@
[…]
@@ 4 Universal Weighted Averaged Stochastic Newton Algorithm
$A_{n}$ $=A_{n-1}-\gamma_{n}\left(P_{n}Q_{n}^{T}+Q_{n}P_{n}^{T}-2\,I_{d}\right)\mathbf{1}_{\left\{\left\|Q_{n}\right\|\left\|Z_{n}\right\|\leq\beta_{n}\right\}}$ (8)
$A_{n,\tau}$ $=(1-\frac{\ln(n+1)^{\tau}}{\sum_{k=0}^{n}\ln(k+1)^{\tau}})A_{n-1,\tau}+\frac{\ln(n+1)^{\tau}}{\sum_{k=0}^{n}\ln(k+1)^{\tau}}A_{n}$ (9)
where $(\nu_{n})_{n\geq 1}$ is a sequence of positive real numbers defined for any $n\geq 1$ by $\nu_{n}=c_{\nu}n^{\nu}$ with $c_{\nu}>0$ and $\nu\in(1/2,1-\beta)$ satisfying $\gamma+\nu>3/2$. In addition, $\tau,\tau^{\prime}\geq 0$. The following theorem gives the consistency of the estimates defined by (6) and (7).
@@
Assume that Assumptions (A1) to (A4) hold. Let $\theta_{n}$ and $\theta_{n,\tau^{\prime}}$ be defined as in (6) and (7). Then,
@@
Assume that Assumptions (A1) to (A4) hold. Then $\theta_{n}$ and $\theta_{n,\tau^{\prime}}$ defined by (6) and (7) satisfy for all $\delta>0$
[…]
@@
[…]
@@ 5 Applications
The simulation section of this paper focuses on evaluating the performance of our novel methods, called Universal Stochastic Newton Algorithm (USNA) and Universal Weighted Averaged Stochastic Newton Algorithms (UWASNA). We begin by analyzing the performance of USNA and UWASNA in the context of logistic regression and the geometric median estimation. In both contexts, it was already feasible to employ second-order algorithms such as the Stochastic Newton Algorithms (SNA) and its Weighted Averaged version (WASNA). These two algorithms use the Riccati formula to recursively compute the inverse of the Hessian estimator. By demonstrating comparable results to SNA and WASNA, we establish USNA and UWASNA as viable alternatives with efficient performance. Additionally, we investigate the applicability of our method in challenging scenarios where using the Riccati formula in SNA is not feasible, particularly in estimating $p$-means and parameters of a spherical distribution. In these two cases, we compare the performances of USNA and UWASNA with the one of the Averaged Stochastic Gradient Descent (ASGD) introduced by Polyak and Juditsky, (1992). Through comprehensive simulations, our goal is to verify that USNA and UWASNA consistently exhibit favourable performances even in these contexts. To close this section, we extend our evaluation to real-world datasets to showcase the practicality of deploying USNA and UWASNA algorithm in applications.
@@ 5.1 Choice of the hyperparameters
    Conditions on $\beta$ and $\gamma$: Condition $\beta<\gamma-1/2$ is only used to apply the Robbins-Siegmund theorem. It serves a theoretical purpose, and we advocate for its removal. On the contrary, condition $\beta_{n+1}\gamma_{n+1}\leq 1/2$ is essential. It ensures the positiveness of the estimate of the inverse of the Hessian. For this reason, in all simulations, we set $\beta_{n+1}=\frac{1}{2}\gamma_{n}^{-1}$.

### v3

@@ 1 Introduction
where $\theta_{0}^{SG}$ is an arbitrarily chosen initial value and $(\nu_{n})_{n\geq 1}$ is a sequence of positive real numbers decreasing towards 0. These algorithms have been extensively studied, with asymptotic results found by Pelletier, (1998, 2000) and non-asymptotic results, such as uniform bounds of the quadratic mean error, presented by Moulines and Bach, (2011); Gadat and Panloup, (2017) to name a few. To ensure asymptotic efficiency, an additional step consists of considering an averaged version of the estimates (Polyak and Juditsky,, 1992).
@@
We thus obtain a universal estimate of the inverse of the Hessian, and thus with reduced calculus time. Note that one could consider using random matrices instead of vectors, and this injection can be linked to Randomized Subspace Newton methods (see, for instance, Gower et al., 2019a ). Additionally, various approaches exist for sketching the Hessian (see Pilanci and Wainwright, (2016, 2017), among others). Notably, in the offline iterative literature, several methods approximate the inverse of the Hessian. One of the most well-known is the BFGS algorithm (Shanno,, 1970), along with its stochastic variants (Schraudolph et al.,, 2007; Byrd et al.,, 2016). Other notable approaches include those proposed by Ye et al., (2017) and Agarwal et al., (2017).
To further enhance convergence rate, we also consider its weighted averaged version, as discussed by Mokkadem and Pelletier, (2011); Boyer and Godichon-Baggioni, (2022). We establish the almost sure rates of convergence for the proposed estimates, after making slight modifications. These results remain true for any consistent estimates $\theta_{n}$. Based on this concept, we introduce a universal recursive Newton algorithm and its weighted averaged version. Additionally, we provide their convergence rates and demonstrate the asymptotic efficiency of the weighted averaged estimates.
@@ 2 Framework
where $g(X,\cdot)$ is a convex function and $X$ is a random vector of ${\mathbb{R}}^{p}$. Furthermore, let us suppose that for all $h\in\mathbb{R}^{d}$, there is a subset $\Omega_{h}\subseteq\mathbb{R}^{p}$ such that for all $x\in\Omega_{h}$, the functional $g(x,.)$ is twice-differentiable at $h$ and $\mathbb{P}\left[X\in\Omega_{h}\right]=1$. We assume also that there exists a unique value $\theta\in{\mathbb{R}}^{d}$ such that
@@
(A1a) There exists $C>0$ such that for all $h\in\mathbb{R}^{d}$,
@@
(A1b) There are positive constants $R,C_{R}>0$ such that for all $h\in\mathbb{R}^{d}$,
$\mathbb{E}\left[\left\|\nabla_{h}g(X,h)\right\|^{4}\mathbf{1}_{\|h-\theta\|\leq R}\right]\leq C_{R}.$
@@
These assumptions are very close to those presented in the literature (Pelletier,, 2000; Gadat and Panloup,, 2017; Godichon-Baggioni,, 2019). Assumption (A1) controls the growth of the gradient and guarantees the stability of the estimation process. It ensures that the gradient remains bounded as the estimation progresses. Observe that Assumption (A1b) plays a crucial role in determining the convergence rate of Newton estimates and is already required to establish the almost sure convergence rate of stochastic gradient algorithms (Pelletier,, 1998). However, it can be relaxed to accommodate any order, depending on the choice of the step size (see, for instance, Boyer and Godichon-Baggioni, (2022); Pelletier, (1998)). For readability, we set it to 4 in this work. Additionally, note that this assumption differs slightly from equation (10) in Boyer and Godichon-Baggioni, (2022), but all results in that paper remain valid under this assumption. Assumption (A2) ensures that the curvature of $G$ at $\theta$ is well-behaved, allowing the estimation algorithm to reliably exploit the local structure of $G$ and is present in most of the offline Newton’s literrature (see Byrd et al., (2016); Ye et al., (2017); Agarwal et al., (2017) among others). This assumption also guarantees that the gradient of $G$ is Lipschitz continuous with a constant $L_{\nabla G}$. This Lipschitz continuity is crucial as convergence results are obtained using a Taylor’s expansion of $G$ up to the second order. Note that Assumption (A1a), when combined with (A2), can be related to the well-known expected smoothness condition (Gower et al., 2019b, ). Assumption (A3) indicates that the Hessian matrix does not exhibit abrupt changes within a neighborhood around $\theta$ and can be related to self-concordance property (Bach,, 2014). This assumption ensures the stability of the Hessian estimates during the estimation process. Assumption (A4) guarantees that the second-order derivative of $G$ does not exhibit excessive fluctuations. It imposes a bound on the variation of the Hessian matrix, providing further stability to the estimation algorithm. It’s worth noting that Hölder’s inequality leads to
@@
These assumptions, along with the differentiability properties of $g$ and $G$, provide a solid foundation for developing efficient second order methods to solve the minimization problem and obtain reliable estimates of $\theta$. Appendix A provides a detailed verification of these assumptions in the contexts of linear and logistic regressions, as well as for estimating the geometric median and $p$-means.
@@ 3 Estimation of the Hessian inverse
In this section, our focus is solely on estimating the inverse Hessian of function $G$ in $\theta$, denoted as $H^{-1}$ with $H=\nabla^{2}G(\theta)$. Even if our motivation is to estimate $H^{-1}$ for proposing a second-order algorithm, the estimation of $H^{-1}$ can also be valuable for recursively constructing confidence intervals or significance statistical tests for a component of parameter $\theta$ when the parameter $\theta$ is estimated using another asymptotically efficient algorithm like the averaged stochastic gradient algorithm. Indeed, in most cases, the asymptotic variance involved in the central limit theorem generally depends on matrix $H^{-1}$ and its estimation is then required.
@@
where $\widehat{A}_{0}$ is an arbitrary symmetric positive definite matrix, and $\gamma_{n}=c_{\gamma}n^{-\gamma}$ with $\frac{1}{2}<\gamma<1$ and $c_{\gamma}>0$. It is important to note that $\widehat{A}_{n}$ is symmetric for any $n\geq 1$ due to its construction. However, since $\theta$ is unknown, we need to estimate it. Assuming we have an efficient recursive estimator $\hat{\theta}_{n-1}$ of $\theta$ (e.g., a stochastic gradient estimator), we can easily derive an estimator of $H^{-1}$ using a plug-in procedure:
@@
This estimator is always symmetric but not necessarily positive definite. To partially ensure positive definiteness, we introduce a truncation based on the norm of $\nabla_{h}^{2}g(X_{n},\widehat{\theta}_{n-1})$, leading to the following estimator of $H^{-1}$:
$\widehat{A}_{n}=\widehat{A}_{n-1}-\gamma_{n}\left(\widehat{A}_{n-1}\nabla_{h}^{2}g(X_{n},\widehat{\theta}_{n-1})+\nabla_{h}^{2}g(X_{n},\widehat{\theta}_{n-1})-2I_{d}\right)\mathbf{1}_{\left\{\left\|\nabla^{2}_{h}g(X_{n},\widehat{\theta}_{n-1})\right\|_{op}\leq\beta_{n}\right\}},$ (3)
where $\beta_{n}=c_{\beta}n^{\beta}$ with $\frac{1-\gamma}{q-1}<\beta<\gamma-\frac{1}{2}$ and $\gamma_{n}\beta_{n}\leq 1/2$. Additionally, this truncation enables control over the smallest eigenvalue of $\widehat{A}_{n}$, which is useful for studying an estimator of the parameter $\theta$ involving the matrix $\widehat{A}_{n}$. This is particularly important in establishing the consistency of the Stochastic Newton algorithm presented in Section 4.
@@
$A_{n}$ $={\Pi_{\beta_{n}^{\prime}}}\left(A_{n-1}-\gamma_{n}\left(P_{n}Q_{n}^{T}+Q_{n}P_{n}^{T}-2\,I_{d}\right)\mathbf{1}_{\{\left\|Q_{n}\right\|\left\|Z_{n}\right\|\leq\beta_{n}\}}\right)$ (4)
where $A_{0}$ is an arbitrary symmetric and positive definite matrix and $\Pi_{\beta_{n}^{\prime}}$ is the projection onto the ball of radius $\beta_{n}^{\prime}$ (with respect to the Frobenius norm), with $\beta_{n}^{\prime}=c_{\beta}^{\prime}n^{\beta^{\prime}}$ satisfying for all $n$, $\gamma_{n}^{2}\beta_{n}^{2}\beta_{n}^{\prime}\leq\gamma_{n}$ and $\beta^{\prime}>1-\gamma$ (i.e. one has to take $\beta^{\prime}\in(1-\gamma,\gamma-2\beta)$, which is possible as soon as $\beta<\gamma-1/2$). Observe that the projection step only necessitates $O(d^{2})$ operations and ensures the largest eigenvalue of $A_{n}$ not to be too large, and also enables to ensure the positive definiteness of $A_{n}$.
We can observe that in this algorithm the truncation is only based on $\left\|Q_{n}\right\|\left\|Z_{n}\right\|$, and not on $\left\|Q_{n}Z_{n}^{T}\right\|_{op}$ as expected, because we have $\left\|Q_{n}Z_{n}^{T}\right\|_{op}=\left\|Q_{n}\right\|\left\|Z_{n}\right\|$. Notably, the computational complexity of each update of $A_{n}$ is now reduced to $\mathcal{O}(d^{2})$. Moreover, following Mokkadem and Pelletier, (2011); Boyer and Godichon-Baggioni, (2022), we can propose a weighted averaged estimate $A_{n,\tau}$, which performs better in practice when the initialization is poor. It is given by
@@
Assume that Assumptions (A2) to (A4) hold, that $\frac{1-\gamma}{q-1}<\beta<\gamma-\frac{1}{2}$, and that there is an estimate $\hat{\theta}_{n}$ satisfying for all $\delta>0$
@@
$\left\|A_{n}-H^{-1}\right\|^{2}=o\left(\frac{\ln n^{1+\delta}}{n^{\min\{\gamma,a,2\beta(q-1)\}}}\right)a.s.\quad\text{and}\quad\left\|A_{n,\tau}-H^{-1}\right\|^{2}=o\left(\frac{\ln n^{1+\delta}}{n^{\min\{1,a,2\beta(q-1)\}}}\right)a.s.$
@@ 4 Universal Weighted Averaged Stochastic Newton Algorithm
$A_{n}$ $={\Pi_{\beta_{n}^{\prime}}}\left(A_{n-1}-\gamma_{n}\left(P_{n}Q_{n}^{T}+Q_{n}P_{n}^{T}-2\,I_{d}\right)\mathbf{1}_{\left\{\left\|Q_{n}\right\|\left\|Z_{n}\right\|\leq\beta_{n}\right\}}\right)$ (8)
$A_{n,\tau}$ $=\left(1-\frac{\ln(n+1)^{\tau}}{\sum_{k=0}^{n}\ln(k+1)^{\tau}}\right)A_{n-1,\tau}+\frac{\ln(n+1)^{\tau}}{\sum_{k=0}^{n}\ln(k+1)^{\tau}}A_{n}$ (9)
where $(\nu_{n})_{n\geq 1}$ is a sequence of positive real numbers defined for any $n\geq 1$ by $\nu_{n}=c_{\nu}n^{-\nu}$ with $c_{\nu}>0$ and $\nu\in(1/2,1-\beta)$ satisfying $\gamma+\nu>3/2$. In addition, $\tau,\tau^{\prime}\geq 0$. The following theorem gives the consistency of the estimates defined by (6) and (7).
@@
Assume that Assumptions (A1a), (A2) to (A4) hold. Suppose also that $\frac{1-\gamma}{q-1}<\beta<\gamma-\frac{1}{2}$, $\gamma+\nu>3/2$, $\nu\in(1/2,1-\beta)$ and $\beta^{\prime}\in(1-\gamma,\gamma-2\beta)$. Let $\theta_{n}$ and $\theta_{n,\tau^{\prime}}$ be defined as in (6) and (7). Then,
@@
Assume that Assumptions (A1) to (A4) hold. Suppose also that $\frac{1-\gamma}{q-1}<\beta<\gamma-\frac{1}{2}$, $\gamma+\nu>3/2$, $\nu\in(1/2,1-\beta)$ and $\beta^{\prime}\in(1-\gamma,\gamma-2\beta)$. Then $\theta_{n}$ and $\theta_{n,\tau^{\prime}}$ defined by (6) and (7) satisfy for all $\delta>0$
$\left\|\theta_{n}-\theta\right\|^{2}=o\left(\frac{\ln n^{1+\delta}}{n^{\nu}}\right)a.s.\quad\text{and}\quad\left\|\theta_{n,\tau^{\prime}}-\theta\right\|^{2}=o\left(\frac{\ln n^{1+\delta}}{n}\right)a.s.$
@@
$\left\|A_{n}-H^{-1}\right\|^{2}=o\left(\frac{\ln n^{1+\delta}}{n^{\gamma}}\right)a.s.\quad\text{and}\quad\left\|A_{n,\tau}-H^{-1}\right\|^{2}=o\left(\frac{\ln n^{1+\delta}}{n}\right)a.s.$
@@ 5 Applications
The simulation section of this paper focuses on evaluating the performance of our novel methods, called Universal Stochastic Newton Algorithm (USNA) and Universal Weighted Averaged Stochastic Newton Algorithm (UWASNA). We begin by analyzing the performance of USNA and UWASNA in the context of logistic regression and the geometric median estimation. In both contexts, it was already feasible to employ second-order algorithms such as the Stochastic Newton Algorithm (SNA) and its Weighted Averaged version (WASNA). These two algorithms use the Riccati formula to recursively compute the inverse of the Hessian estimator. By demonstrating comparable results to SNA and WASNA, we establish USNA and UWASNA as viable alternatives with efficient performance. Additionally, we investigate the applicability of our method in challenging scenarios where using the Riccati formula in SNA is not feasible, particularly in estimating $p$-means and parameters of a spherical distribution. In these two cases, we compare the performances of USNA and UWASNA with the one of the Averaged Stochastic Gradient Descent (ASGD) introduced by Polyak and Juditsky, (1992). Through comprehensive simulations, our goal is to verify that USNA and UWASNA consistently exhibit favourable performances even in these contexts. To close this section, we extend our evaluation to real-world datasets to showcase the practicality of deploying USNA and UWASNA algorithm in applications.
@@ 5.1 Choice of the hyperparameters
    Conditions on $\beta$ and $\gamma$: In practice, we find that the condition $\beta<\gamma-1/2$ is not necessary for the convergences of the estimates, and we advocate for its removal. On the contrary, condition $\beta_{n+1}\gamma_{n+1}\leq 1/2$ is essential. It ensures the positiveness of the estimate of the inverse of the Hessian. For this reason, in all simulations, we set $\beta_{n}=\frac{1}{2}n^{3/4}$ and $\gamma_{n}=n^{-3/4}$.